# Optimizing quantum battery performance by reducing battery influence in charging dynamics

### v1

@@ III.2 Interacting battery and non-interacting charger Hamiltonian
Figure 4: (a) $\Delta E$ and (b) $P$ vs. $t$ for NN Ising interacting Hamiltonian as a battery and non-interacting Hamiltonian as a charger. (c) $\Delta E$ and (d) $P$ vs. $t$ for NN interacting XY spin system as a battery and non-interacting Hamiltonian as a charger. Parameterse are: $N=12$, $J=1$, $h=1$, and $\gamma=0.5$. Periodic boundary conditions are considered.
@@
We analyze the storage energy and power of the battery in two scenarios: one where the battery’s influence on the charger is present ($\lambda=0$) and another where it is absent ($\lambda=1$). The results demonstrate that, for both Ising and XY interacting batteries, storage energy and power are notably higher when $\lambda=1$, underscoring the beneficial role of the battery’s countereffect in the charging dynamics. Additionally, the Ising model consistently yields superior performance over the XY model in both metrics, as shown in Fig. 4(a–d).
@@ III.3 Interacting battery and interacting charger Hamiltonian
We consider the scenario where both the battery and charger are governed by interacting Hamiltonians. First, we analyze the case where the battery is modeled by an Ising spin system with NN interactions [Eq. (10)], and the charger is described by an NN XY-interacting Hamiltonian [Eq. (12)]. We calculate the storage energy and power in two situations: when the battery’s influence is present in the charging dynamics ($\lambda=0$) and when it is absent ($\lambda=1$). Our results indicate that the presence of the counter effect of battery during charging leads to an advantage in both the stored energy and the power [Fig. 5(a,b)].
@@
Figure 5: (a) $\Delta E$ and (b) $P$ as a function of time $t$ for a NN interacting Ising spin system as the battery and NN interacting XY spin system as the charger Hamiltonian. (c) $\Delta E$ and (d) $P$ vs. $t$ for a NN interacting XY spin system as the battery and NN interacting Ising spin system as the charger Hamiltonian. The system parameters are $N=12$, $J=1$, and $\gamma=0.5$. Periodic boundary conditions are considered.
@@ IV Conclusion
In this work, we explored the role of the battery’s countereffect in the charging dynamics of quantum batteries composed of spin systems. By analyzing various configurations involving both non-interacting and interacting Hamiltonians for the battery and charger, we demonstrated how their interplay significantly impacts storage energy and power.
We began by examining a non-interacting battery coupled to an interacting charger modeled by an ATA Ising spin system. Our results revealed that incorporating the battery’s countereffect into the dynamics ($\lambda=1$) leads to optimal performance, with the storage energy reaching its theoretical maximum $2Nh$ and the power also peaking. Notably, an odd-even parity effect emerged in energy storage for increasing system sizes, while power remained parity-independent.
To further understand the influence of interaction types and ranges, we compared Ising and XY charger models with both ATA and NN interactions. Across all these cases, the optimum storage energy and power consistently occurred at $\lambda=1$, emphasizing the constructive role of the countereffect in enhancing energy transfer. Both quantities exhibited linear scaling with system size, confirming the scalability of our protocol.
Extending the analysis to interacting batteries with non-interacting chargers, where the battery is governed by either Ising or XY NN interactions, we observed that the inclusion of the countereffect continued to offer superior energy and power outcomes. This advantage persisted even in the most general scenario where both the battery and charger are interacting systems, reinforcing the robustness of our findings across diverse interaction frameworks.
Interestingly, in cases where the charger is modeled with ATA interactions, the charging power peaks beyond the physically meaningful range ($\lambda>1$). Although this shift lies outside the regime of practical relevance, it suggests the presence of underlying physical mechanisms that merit deeper investigation. A detailed discussion of this anomaly is provided in Appendix A.
Altogether, our study establishes the critical importance of retaining the battery’s countereffect in the charger Hamiltonian for optimizing quantum battery performance. These insights offer valuable guidance for the design of efficient and scalable quantum energy storage devices and open new avenues for exploring interaction-driven enhancements in quantum thermodynamic protocols.
@@ Appendix A Charging dynamics beyond physical regime
Figure 6: Storage power $P(t)$ as a function of time $t$ for ATA interacting (a) Ising and (b) $XY$ charger Hamiltonian, shown for various values of the countereffect parameter $\lambda$. The insets display the corresponding maximum storage power as a function of $\lambda$. Parameters used: $J=h=1$, $N=10$, and $\gamma=0.5$.

### v2

@@ III.2 Interacting battery and non-interacting charger Hamiltonian
Figure 4: (a) Stored energy $\Delta E$ and (b) charging power $P$ as functions of time $t$ for a battery described by a NN Ising Hamiltonian with a non-interacting charger. (c) Stored energy $\Delta E$ and (d) charging power $P$ for a battery described by a NN XY Hamiltonian with the same non-interacting charger. The system parameters are $N=12$, $J=1$, $h=1$, and $\gamma=0.5$, with periodic boundary conditions.
@@
Within this model, the quantum battery is modeled as an interacting chain of qubits whose internal spin–spin couplings facilitate collective energy storage and the buildup of correlations during the charging process. When the charging protocol is initiated at time $t_{\mathrm{on}}$, a noninteracting charger—realized as a uniform external field—injects energy coherently into all qubits without generating additional correlations. To account for the feedback-like countereffect during charging, the internal interaction strength of the battery is simultaneously modulated as $J\rightarrow(1-\lambda)J$, where $\lambda$ quantifies the degree of suppression of the battery’s intrinsic contribution. This configuration, representing an interacting battery driven by a noninteracting charger, provides a clear framework to explore how intrinsic interactions influence the redistribution of injected energy, thereby enhancing storage capacity, charging power, and the overall robustness of the quantum battery.
We analyze the storage energy and power of the battery in two scenarios: one where the battery’s countereffec on the charger is absent ($\lambda=0$) and another where it is present ($\lambda=1$). The results demonstrate that, for both Ising and XY interacting batteries, storage energy and power are notably higher when battery counter effect is maximal in the charger ($\lambda=1$), underscoring the beneficial role of the battery’s countereffect in the charging dynamics. Additionally, the Ising model consistently yields superior performance over the XY model in both metrics, as shown in Fig. 4(a–d).
We investigate how the storage energy and power depend on the interaction strength of the battery. In this analysis, the countereffect of the battery on the charger is not included, so as to clearly isolate the influence of interactions on storage energy and power. Our results show that the storage energy increases with $J$, reaches a maximum at a particular time, and then decreases, while the time required to attain the maximum also decreases as the interaction strength $J$ increases (Fig. 5(a)). The inset of Fig. 5(a) highlights how the maximum storage energy varies with $J$: it grows as $J$ approaches $1$, reaches its peak at $J=1$, and then decreases beyond this point. Notably, the critical value $J=1$ coincides with the charger strength $h=1$, suggesting that optimal storage energy is achieved when the battery and charger strengths are equal. The power follows a similar trend: it grows with time, reaches a maximum, and then decays for all values of $J$. Moreover, the time to reach the maximum power decreases as $J$ increases (Fig. 5(b)). The maximum storage power exhibits a logarithmic dependence on $J$, approximately following $P_{\rm max}\approx 17.91\log_{10}(J)+16.86$ (inset of Fig. 5(b)). Similar observations were reported in Ref. [ghosh2020enhancement], which demonstrated that both ordered and disordered interaction strengths significantly enhance power extraction. Additionally, Ref. [PhysRevA.105.022628] showed that the average power output of a quantum battery based on an interacting spin system, charged via a local magnetic field, increases with the spin quantum number, highlighting a dimensional advantage in quantum batteries.
Figure 5: Storage energy $\Delta E$ and power $P$ of the NN Ising spin system as a quantum battery, with non-interacting spins as the charger, shown as a function of time for different interaction strengths $J$ (fixing $\lambda=0$). The inset shows the maximum storage energy and power as $J$ is varied. The maximum power scales approximately logarithmically with interaction strengh , following $P_{\rm max}\approx 17.91\log_{10}(J)+16.86$.
@@ III.3 Interacting battery and interacting charger Hamiltonian
Figure 6: (a) Stored energy $\Delta E$ and (b) charging power $P$ as functions of time $t$ for a battery described by a NN Ising spin system and a charger described by a NN XY spin system. (c) Stored energy $\Delta E$ and (d) charging power $P$ for the case where the battery is a NN XY spin system and the charger is a NN Ising spin system. The calculations use $N=12$, $J=1$, and $\gamma=0.5$ with periodic boundary conditions.
We consider the scenario in which both the battery and the charger are interacting spin systems. In the first case, the battery is modeled as an Ising spin chain with NN interactions [Eq. (8)], while the charger is implemented as a nearest-neighbor XY-interacting Hamiltonian [Eq. (10)]. In the reversed configuration, the roles are swapped: the battery is described by a NN XY-interacting chain, and the charger is realized as an NN Ising spin system.
In this configuration, the system is initially prepared in the ground state of an Ising-type Hamiltonian, representing an interacting quantum battery characterized by spin–spin couplings that enable collective energy storage. At the onset of charging ($t=t_{\mathrm{on}}$), the interaction profile of the system is suddenly modified by switching on an XY-type Hamiltonian, which now acts as an effective charger. This sudden quench introduces flip-flop interactions that coherently inject energy into the spin system, promoting excitation delocalization and the build-up of correlations among the qubits. The two Hamiltonians—Ising and XY—thus represent distinct dynamical regimes of the same physical platform, allowing the same spin chain to serve sequentially as both the battery and the charger. To investigate the role of the countereffect, the intrinsic Ising interaction strength is simultaneously reduced as $J\rightarrow(1-\lambda)J$ during charging, where $\lambda$ controls the degree to which internal correlations are suppressed. This setup provides a unified framework to explore how the interplay between Ising and XY interactions governs energy absorption, redistribution, and storage efficiency, and how the externally applied countereffect modifies these cooperative charging dynamics.
We calculate the storage energy and power in two situations: when the battery’s influence is present in the charging dynamics ($\lambda=0$) and when it is absent ($\lambda=1$). Our results indicate that the presence of the counter effect of battery during charging leads to an advantage in both the stored energy and the power [Fig. 6(a,b)].
@@ IV Conclusion
This work presents a comprehensive study of charging dynamics in spin-based quantum batteries, focusing on the often-overlooked influence of the battery’s countereffect during the charging dynamics. By systematically analyzing configurations involving interacting and non-interacting Hamiltonians for both the battery and the charger, we reveal a consistent and compelling theme: retaining the battery’s countereffect significantly boosts charging performance.
Our findings show that maximum storage energy and power are achieved when the countereffect is fully included ($\lambda=1$), regardless of whether the system features ATA or NN interactions, and whether the interactions are of the Ising or XY type. Particularly, in the case of a non-interacting battery with an ATA Ising charger, the stored energy reaches its upper bound $2Nh$, accompanied by a notable odd-even parity effect in energy scaling—though power remains unaffected by parity.
Importantly, this enhancement remains robust when the roles are reversed, i.e., with an interacting battery and a non-interacting charger. Even in the most general case, where both battery and charger are interacting many-body systems, the countereffect continues to improve energy transfer and power, suggesting it as a universal design principle for efficient quantum batteries.
In the case of a NN Ising interaction battery coupled to a non-interacting charger, the battery’s interaction strength enhances the maximum storage energy up to the point where the interaction strength matches that of the charger. However, the maximum power increases approximately logarithmically with the interaction strength.
A curious feature emerges in some ATA models where peak power occurs at $\lambda>1$, outside the regime of physical interpretability. While not directly useful for device implementation, this behavior indicates nontrivial underlying dynamics, which could inspire future theoretical investigations (see Appendix A).
Altogether, our results uncover a simple yet powerful insight: the battery’s back-action on the charger is not a hindrance but a resource. Incorporating this countereffect in the system design leads to substantial gains in energy and power storage, offering valuable direction for next-generation quantum energy storage technologies.
@@ Appendix A Charging dynamics beyond physical regime
Figure 7: Charging power $P(t)$ as a function of time $t$ for a non-interacting battery charged by an ATA interacting (a) Ising and (b) XY Hamiltonian. Results are shown for several values of the countereffect parameter $\lambda$. The insets display the corresponding maximum charging power as a function of $\lambda$. The simulations use $J=1$, $h=1$, system size $N=10$, and anisotropy parameter $\gamma=0.5$.